A constant force 𝑮⃗ ⃗

= (𝟓. 𝟏𝟏𝒋̂ + 𝟔. 𝟏𝟏 𝒌 ̂)𝑵 is applied to a particle that

undergoes a displacement 𝓵⃗ ⃗ ⃗

= (𝟒. 𝟏𝟏𝒋̂  − 𝟕. 𝟏𝟏 𝒌 ̂)𝒏. (a) the work done by 

the force, and (b) the component of the force in the direction of the 

displacement.


1
Expert's answer
2021-10-26T09:42:00-0400

(a) The work is


"W=\vec G·\vec l=5.11·4.11+6.11·(-7.11)=
27.1\text { J}."

(b) The component of the force is


"G_l=W/|l|=27.1/\sqrt{4.11^2+7.11^2}=3.30\text{ N}."